OpenStudy (anonymous):

John gave the following explanation for factoring the trinomial x2 - x - 2: "I need to find factors of -2. -1 and 2 multiply to -2 so my factors are (x - 1)(x + 2)."

OpenStudy (anonymous):

Part 1: Are John's factors correct? If not, explain where he made his mistake. Part 2: What could John have done to check the accuracy of his factors before submitting his Module 3 test?

OpenStudy (anonymous):

@amistre64 could you help me with this one too?

OpenStudy (marissalovescats):

Johns factors are not correct. His two binomials of (x-1)(x+2) would come out to make: x times x = x^2 2x-1x=x -1 times 2 is -2 x^2+x-2 is not x^2-x-2. So what do you think we can do to fix that?

OpenStudy (mathstudent55):

Part 1. His factors are incorrect. To factor that polynomial you do need two factors of -2, but they need to add to -1. In this case, the two factors he chose, -1 and 2 do not add to -1, so they are not the correct factors.

OpenStudy (marissalovescats):

To check your factors to ensure they are correct, you do what I just did. You FOIL your two binomials to see if it comes out to the trinomial you started with.

OpenStudy (mathstudent55):

Part 2. He could have multiplied out his factored form to see if he ends up with the original polynomial. If he doesn't, then he knows the factoring is wrong.